OpenStudy (anonymous):

A stone is thrown upward from ground level. The initial speed is 160 feet per second. (a) In how many seconds will the stone hit the ground?

OpenStudy (anonymous):

does not match with my answer choices

OpenStudy (anonymous):

lets calculate it

OpenStudy (anonymous):

the answers i am given are 16, 14,10,9,12

OpenStudy (anonymous):

a) i'm just gonna calculate the time it takes to reach the peak, and multiply that time by two, because the time going up will equal the time going down. (v0 = initial velocity BTW) v will be set to zero, because the only time velocity is 0 is when the projectile reaches its peak, which is what we want v = v0 + at 0 = 160 + -32t -160 = -32t t = 5s takes 5s to go up. multiply by two, it will hit the ground in 10 seconds

OpenStudy (anonymous):

This is a physics problem. Post it in the Physics section and I'll be glad to help. Also note that you'll need your kinematics equations for this. @megumi

OpenStudy (anonymous):

Answer: 10 seconds

OpenStudy (anonymous):

10

OpenStudy (anonymous):

it is a Calculus problme

OpenStudy (anonymous):